OpenStudy (anonymous):

I have a problem here a^2 + 8a +13 =0 Now I have the answer but I dont understand how they got there, I thought well ok this looks like something I could factor out but I realized you cannot factor it but you can use the quadratic equation and I tried but I got a stupid answer, -3 radical 3 or -5 radical 3

OpenStudy (anonymous):

your answer should be -4+sqrt(3) and -4-sqrt(3)

OpenStudy (anonymous):

thank you irene, however I had the answer already I just dont know how to get there.

OpenStudy (anonymous):

on what grade are you now?

OpenStudy (anonymous):

its the first course of college math for me I am a freshman. Its called college algebra withe trig

OpenStudy (anonymous):

just use quadratic equation

OpenStudy (unklerhaukus):

\[ a^2 + 8a +13 =0\] \[(\color{red}1)a^2 + (\color{blue}8)a +(\color{green}{13}) =0\] \[a_{1,2}=\frac{-(\color{blue}8)\pm\sqrt{(\color{blue}8)^2-4(\color{red}1)(\color{green}{13})}}{2\times(\color{red}1)}\]\[\qquad=\quad \dots\]

OpenStudy (unklerhaukus):

\[\qquad=\frac{-8\pm\sqrt{12}}{2}\]\[\qquad=\frac{-8\pm\sqrt{2^2\times3}}{2}\]\[\qquad=\frac{-4\times2\pm2\sqrt{3}}{2}\] \[\qquad=-4\pm\sqrt{3}\]